# Existence and Uniqueness of Solutions for Steady Triple-Deck Equations

### v1

@@ 2. Preliminaries
Fixing $\alpha>\frac{2}{3},\ \varepsilon\in(0,\frac{1}{6}]$, suppose $\omega\in X_{0}\cap X_{\alpha}$ and $\partial_{x}A_{M}\in L^{\infty}\cap H^{\frac{\alpha}{2}}$, then we have $h\in H_{x}^{\frac{\alpha}{2}}L_{y}^{2}$. Moreover, there holds
(2.12) ${\big{\|}h\big{\|}}_{L^{2}(\Omega_{M})}\leq CM^{\frac{1+3\varepsilon}{2}}\bigg{(}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{2}({\mathbb{R}})}\bigg{)}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}},\quad\forall\ \varepsilon\in(0,\frac{1}{6}]$
${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}h\big{\|}}_{L^{2}(\Omega_{M})}\leq CM\bigg{(}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{\infty}({\mathbb{R}})\cap\dot{H}^{\frac{\alpha}{2}}({\mathbb{R}})}\bigg{)}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}.$
Here $C=C(\alpha,\varepsilon)>0$ is independent of $M,\ \omega$, and $A_{M}$.
@@
In view of Proposition 2.1 and Proposition 2.2, it is straight to obtain
@@
$\leq$ $\ C({\big{\|}\omega\big{\|}}_{L^{\infty}(\Omega_{M})}+M^{\frac{3\varepsilon}{2}}{\big{\|}y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}){\big{\|}y\partial_{x}\omega\big{\|}}_{L^{2}(\Omega_{M})}$
@@
$\leq$ $\ C(1+M^{\frac{1+3\varepsilon}{2}}){\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+CM^{\frac{1+3\varepsilon}{2}}{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{2}({\mathbb{R}})}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}.$
which implies the first desired result.
For the high-order estimates, by product estimates, we have
${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}(u\partial_{x}\omega)\big{\|}}_{L^{2}(\Omega_{M})}$
$\qquad\leq C{\big{\|}\omega\big{\|}}_{L^{\infty}(\Omega_{M})}{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y\partial_{x}\omega\big{\|}}_{L^{2}(\Omega_{M})}+C{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\omega\big{\|}}_{H_{x}^{\frac{2}{3}}L_{y}^{2}}{\big{\|}y|\partial_{x}|^{\frac{1}{3}}\omega\big{\|}}_{L^{\infty}(\Omega_{M})}$
$\qquad\leq C{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+C{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\omega\big{\|}}_{H_{x}^{\frac{2}{3}}L_{y}^{2}}{\big{\|}y|\partial_{x}|^{\frac{1}{3}}\omega\big{\|}}_{L^{\infty}(\Omega_{M})},$
@@
${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}(V^{0}\partial_{y}\omega)\big{\|}}_{L^{2}(\Omega_{M})}$
$\qquad\leq C{\big{\|}y^{-\frac{1}{2}}\int_{0}^{y}\int_{z}^{M}|\partial_{x}|^{\frac{\alpha}{2}}\partial_{x}\omega(\cdot,\tilde{z})d\tilde{z}dz\big{\|}}_{L_{x}^{2}L_{y}^{\infty}}{\big{\|}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{\infty}L_{y}^{2}}$
$\qquad\qquad+C{\big{\|}y^{-\frac{1}{2}}\int_{0}^{y}\int_{z}^{M}\partial_{x}\omega(\cdot,\tilde{z})d\tilde{z}dz\big{\|}}_{L_{x}^{4}L_{y}^{\infty}}{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{4}L_{y}^{2}}$
$\qquad\leq CM^{\frac{3\varepsilon}{2}}{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\tilde{z}\partial_{x}\omega\big{\|}}_{L^{2}(\Omega_{M})}{\big{\|}y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}$
$\qquad\qquad+C{\big{\|}\tilde{z}\partial_{x}\omega\big{\|}}_{L_{x}^{4}L_{\tilde{z}}^{2}}{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{4}L_{y}^{2}}$
$\qquad\leq CM^{\frac{3\varepsilon}{2}}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+C{\big{\|}\tilde{z}\partial_{x}\omega\big{\|}}_{L_{x}^{4}L_{\tilde{z}}^{2}}{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{4}L_{y}^{2}}.$
All we left is to give the estimates of the right hand of above inequality. By Sobolev embedding Theorem, we have
${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\omega\big{\|}}_{H_{x}^{\frac{2}{3}}L_{y}^{2}}\leq C{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}+C{\big{\|}|\partial_{x}|^{\frac{2}{3}+\frac{\alpha}{2}}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}$
@@
${\big{\|}y\partial_{x}\omega\big{\|}}_{L_{x}^{4}L_{y}^{2}}\leq C{\big{\|}|\partial_{x}|^{\frac{1}{4}}y\partial_{x}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}$
$\qquad\leq C\big{(}{\big{\|}y\partial_{x}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y\partial_{x}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}\big{)}\quad\mbox{for}\ \ \alpha\geq\frac{1}{2}$
$\qquad\leq C{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}},$
@@
$\qquad\leq C\big{(}{\big{\|}|\partial_{x}|^{\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}+\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}\big{)}$
$\qquad\leq C{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}.$
Combining all above estimates, for $\alpha>\frac{2}{3}$, we obtain
${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}(u\partial_{x}\omega)\big{\|}}_{L^{2}(\Omega_{M})}\leq$ $\ CM{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2},$
@@
$\qquad\leq C{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\partial_{x}A_{M}\big{\|}}_{L^{2}({\mathbb{R}})}{\big{\|}y\partial_{y}\omega\big{\|}}_{L_{x}^{\infty}L_{y}^{2}}+C{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{\infty}({\mathbb{R}})}{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y\partial_{y}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}$
$\qquad\leq CM^{\frac{1+3\varepsilon}{2}}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}{\big{\|}\partial_{x}A_{M}\big{\|}}_{\dot{H}^{\frac{\alpha}{2}}}+CM{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{\infty}}$
$\qquad\leq CM{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}.$
@@
(2.14) ${\big{\|}h_{1}\big{\|}}_{L^{2}(\Omega_{M})}\leq CM^{\frac{1+3\varepsilon}{2}}\bigg{(}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{2}({\mathbb{R}})}\bigg{)}{\big{\|}\omega_{1}\big{\|}}_{X_{0}\cap X_{\alpha}},\ \ \forall\varepsilon\in(0,\frac{1}{6}]$
${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}h_{1}\big{\|}}_{L^{2}(\Omega_{M})}\leq CM\bigg{(}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{\infty}({\mathbb{R}})\cap\dot{H}^{\frac{\alpha}{2}}({\mathbb{R}})}\bigg{)}{\big{\|}\omega_{1}\big{\|}}_{X_{0}\cap X_{\alpha}}.$
Next we give priori estimate of $I_{M}[f]$ for $f\in X_{\bar{\alpha}}$.
Proposition 2.4.
Suppose $\bar{\alpha}\geq 0$, for any $f\in X_{\bar{\alpha}}$, we have
@@
Here $C$ is independent of $\bar{\alpha}$ and $f$.
@@
$\big{|}\int_{0}^{M}\hat{f}(\xi,y)dy\big{|}=$ $\ \big{|}\int_{0}^{|\xi|^{-\frac{1}{3}}}\hat{f}(\xi,y)dy+\int_{|\xi|^{-\frac{1}{3}}}^{M}y^{-1}y\hat{f}(\xi,y)dy\big{|}$
@@
which implies that
(2.16) $\big{|}\xi^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}I_{M}[f]\big{|}\leq$ $\ C|\xi|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\cdot{\big{\|}\hat{f}(\xi,\cdot)\big{\|}}_{L_{y}^{2}({\mathbb{R}})}+C|\xi|^{1+\frac{\bar{\alpha}}{2}}{\big{\|}y\hat{f}(\xi,\cdot)\big{\|}}_{L_{y}^{2}}).$
Thus,
(2.17) ${\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}I_{M}[f]\big{\|}}_{L^{2}({\mathbb{R}})}\leq C\big{(}{\big{\|}|\partial_{x}|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}f\big{\|}}_{L^{2}}+{\big{\|}y|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{x}f\big{\|}}_{L^{2}}\big{)}\leq C{\big{\|}f\big{\|}}_{Y_{\bar{\alpha}}},$
which implies the desired results. ∎
3. Proof of Theorem 1.1
In this section, we give the proof of the Theorem 1.1. Before that, we study a nonhomogeneous linearized problem without the integral condition as follows.

### v2

@@ 2. Functional spaces and nonlinear estimates
Let $\alpha>\frac{2}{3},\ \varepsilon\in(0,\frac{1}{6}]$. Assume that $\omega\in X_{0}\cap X_{\alpha}$ and $\partial_{x}A_{M}\in L^{\infty}\cap H^{\frac{\alpha}{2}}$. Then $h\in H_{x}^{\frac{\alpha}{2}}L_{y}^{2}$ with
(2.2) ${\big\|h\big\|}_{L^{2}(\Omega_{M})}\leq CM^{\frac{1+3\varepsilon}{2}}\big({\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\partial_{x}A_{M}\big\|}_{L^{2}({\mathbb{R}})}\big){\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}},$
${\big\||\partial_{x}|^{\frac{\alpha}{2}}h\big\|}_{L^{2}(\Omega_{M})}\leq CM\big({\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\partial_{x}A_{M}\big\|}_{L^{\infty}({\mathbb{R}})\cap\dot{H}^{\frac{\alpha}{2}}({\mathbb{R}})}\big){\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}.$
Here $C=C(\alpha,\varepsilon)>0$ is independent of $M$ and $A_{M}$.
@@
By virtue of Lemma 2.1 and Lemma 2.2, we obtain
@@
$\leq$ $\ C\big({\big\|\omega\big\|}_{L^{\infty}(\Omega_{M})}+M^{\frac{3\varepsilon}{2}}{\big\|y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big\|}_{L_{y}^{2}L_{x}^{\infty}}\big){\big\|y\partial_{x}\omega\big\|}_{L^{2}(\Omega_{M})}$
@@
$\leq$ $\ C(1+M^{\frac{1+3\varepsilon}{2}}){\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}^{2}+CM^{\frac{1+3\varepsilon}{2}}{\big\|\partial_{x}A_{M}\big\|}_{L^{2}({\mathbb{R}})}{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}},$
which gives the first estimate of (2.2).
For the higher-order estimates, by product estimates, we have
(2.3) ${\big\||\partial_{x}|^{\frac{\alpha}{2}}(u\partial_{x}\omega)\big\|}_{L^{2}(\Omega_{M})}$
$\leq C{\big\|\omega\big\|}_{L^{\infty}(\Omega_{M})}{\big\||\partial_{x}|^{\frac{\alpha}{2}}y\partial_{x}\omega\big\|}_{L^{2}(\Omega_{M})}+C{\big\||\partial_{x}|^{\frac{\alpha}{2}}\omega\big\|}_{H_{x}^{\frac{2}{3}}L_{y}^{2}}{\big\|y|\partial_{x}|^{\frac{1}{3}}\omega\big\|}_{L^{\infty}(\Omega_{M})}$
$\leq C{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}^{2}+C{\big\||\partial_{x}|^{\frac{\alpha}{2}}\omega\big\|}_{H_{x}^{\frac{2}{3}}L_{y}^{2}}{\big\|y|\partial_{x}|^{\frac{1}{3}}\omega\big\|}_{L^{\infty}(\Omega_{M})},$
@@
(2.4) ${\big\||\partial_{x}|^{\frac{\alpha}{2}}(V^{0}\partial_{y}\omega)\big\|}_{L^{2}(\Omega_{M})}$
$\leq C{\big\|y^{-\frac{1}{2}}\int_{0}^{y}\int_{z}^{M}|\partial_{x}|^{\frac{\alpha}{2}}\partial_{x}\omega(\cdot,\tilde{z})d\tilde{z}dz\big\|}_{L_{x}^{2}L_{y}^{\infty}}{\big\|y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{\infty}L_{y}^{2}}$
$\qquad+C{\big\|y^{-\frac{1}{2}}\int_{0}^{y}\int_{z}^{M}\partial_{x}\omega(\cdot,\tilde{z})d\tilde{z}dz\big\|}_{L_{x}^{4}L_{y}^{\infty}}{\big\||\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{4}L_{y}^{2}}$
$\leq CM^{\frac{3\varepsilon}{2}}{\big\||\partial_{x}|^{\frac{\alpha}{2}}\tilde{z}\partial_{x}\omega\big\|}_{L^{2}(\Omega_{M})}{\big\|y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big\|}_{L_{y}^{2}L_{x}^{\infty}}$
$\qquad+C{\big\|\tilde{z}\partial_{x}\omega\big\|}_{L_{x}^{4}L_{\tilde{z}}^{2}}{\big\||\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{4}L_{y}^{2}}$
$\leq CM^{\frac{3\varepsilon}{2}}{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}^{2}+C{\big\|\tilde{z}\partial_{x}\omega\big\|}_{L_{x}^{4}L_{\tilde{z}}^{2}}{\big\||\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{4}L_{y}^{2}}.$
It remains to estimate the right-hand side of the above inequality. By the Sobolev embedding, we have
(2.5) ${\big\||\partial_{x}|^{\frac{\alpha}{2}}\omega\big\|}_{H_{x}^{\frac{2}{3}}L_{y}^{2}}\leq C{\big\||\partial_{x}|^{\frac{\alpha}{2}}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}+C{\big\||\partial_{x}|^{\frac{2}{3}+\frac{\alpha}{2}}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}$
@@
(2.6) ${\big\|y\partial_{x}\omega\big\|}_{L_{x}^{4}L_{y}^{2}}$ $\leq C{\big\||\partial_{x}|^{\frac{1}{4}}y\partial_{x}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}$
$\leq C\big({\big\|y\partial_{x}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}+{\big\||\partial_{x}|^{\frac{\alpha}{2}}y\partial_{x}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}\big)\quad\mbox{for}\ \ \alpha\geq\frac{1}{2}$
$\leq C{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}},$
@@
$\leq C\big({\big\||\partial_{x}|^{\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}+{\big\||\partial_{x}|^{\frac{\alpha}{2}+\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}\big)$
$\leq C{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}.$
Combining all estimates above, we deduce that for $\alpha>\frac{2}{3}$,
(2.7) ${\big\||\partial_{x}|^{\frac{\alpha}{2}}(u\partial_{x}\omega)\big\|}_{L^{2}(\Omega_{M})}\leq\ CM{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}^{2},$
@@
$\leq C{\big\||\partial_{x}|^{\frac{\alpha}{2}}\partial_{x}A_{M}\big\|}_{L^{2}({\mathbb{R}})}{\big\|y\partial_{y}\omega\big\|}_{L_{x}^{\infty}L_{y}^{2}}+C{\big\|\partial_{x}A_{M}\big\|}_{L^{\infty}({\mathbb{R}})}{\big\||\partial_{x}|^{\frac{\alpha}{2}}y\partial_{y}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}$
$\leq CM^{\frac{1+3\varepsilon}{2}}{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}{\big\|\partial_{x}A_{M}\big\|}_{\dot{H}^{\frac{\alpha}{2}}}+CM{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}{\big\|\partial_{x}A_{M}\big\|}_{L^{\infty}}$
$\leq CM{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}{\big\|\partial_{x}A_{M}\big\|}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}.$
Substituting (2.5)-(2.8) into (2.3)-(2.4), we conclude the last estimate of (2.2). ∎
@@
${\big\|h_{1}\big\|}_{L^{2}(\Omega_{M})}\leq CM^{\frac{1+3\varepsilon}{2}}\big({\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\partial_{x}A_{M}\big\|}_{L^{2}({\mathbb{R}})}\big){\big\|\omega_{1}\big\|}_{X_{0}\cap X_{\alpha}},\ \ \varepsilon\in(0,\frac{1}{6}],$
${\big\||\partial_{x}|^{\frac{\alpha}{2}}h_{1}\big\|}_{L^{2}(\Omega_{M})}\leq CM\big({\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\partial_{x}A_{M}\big\|}_{L^{\infty}({\mathbb{R}})\cap\dot{H}^{\frac{\alpha}{2}}({\mathbb{R}})}\big){\big\|\omega_{1}\big\|}_{X_{0}\cap X_{\alpha}}.$
The following lemma is devoted to the integral boundary condition $I_{M}[\omega]=A(x)+F(x)$.
Lemma 2.3.
It holds that for any $\bar{\alpha}\geq 0$,
@@
Here $C$ is independent of $\bar{\alpha}$.
@@
$\left|\int_{0}^{M}\hat{f}(\xi,y)dy\right|=$ $\left|\int_{0}^{|\xi|^{-\frac{1}{3}}}\hat{f}(\xi,y)dy+\int_{|\xi|^{-\frac{1}{3}}}^{M}y^{-1}y\hat{f}(\xi,y)dy\right|$
@@
which implies
$\big|\xi^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}I_{M}[f]\big|\leq$ $\ C|\xi|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\cdot{\big\|\hat{f}(\xi,\cdot)\big\|}_{L_{y}^{2}({\mathbb{R}})}+C|\xi|^{1+\frac{\bar{\alpha}}{2}}{\big\|y\hat{f}(\xi,\cdot)\big\|}_{L_{y}^{2}}).$
Thus, we obtain
${\big\||\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}I_{M}[f]\big\|}_{L^{2}({\mathbb{R}})}\leq C\big({\big\||\partial_{x}|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}f\big\|}_{L^{2}}+{\big\|y|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{x}f\big\|}_{L^{2}}\big)\leq C{\big\|f\big\|}_{Y_{\bar{\alpha}}}.$
It completes the proof. ∎
3. The solvability of the linear system
Recall that, to solve the nonlinear system (1.13), we decompose $\omega=\omega_{0}+\bar{\omega}$ and $A=A_{0}+\bar{A}$ with
and
To address the system (3.2), we first solve an inhomogeneous linear problem without the integral condition